# NatureLM: Deciphering the Language of Nature for Scientific Discovery

### v1

@@ Abstract
Foundation models have revolutionized natural language processing and artificial intelligence, significantly enhancing how machines comprehend and generate human languages. Inspired by the success of these foundation models, researchers have developed foundation models for individual scientific domains, including small molecules, materials, proteins, DNA, and RNA. However, these models are typically trained in isolation, lacking the ability to integrate across different scientific domains. Recognizing that entities within these domains can all be represented as sequences, which together form the “language of nature”, we introduce Nature Language Model (briefly, NatureLM), a sequence-based science foundation model designed for scientific discovery. Pre-trained with data from multiple scientific domains, NatureLM offers a unified, versatile model that enables various applications including: (i) generating and optimizing small molecules, proteins, RNA, and materials using text instructions; (ii) cross-domain generation/design, such as protein-to-molecule and protein-to-RNA generation; and (iii) achieving state-of-the-art performance in tasks like SMILES-to-IUPAC translation and retrosynthesis on USPTO-50k.
@@ 1 Introduction
Despite their success in general-purpose tasks, early investigations ai4science2023impactlargelanguagemodels highlight significant room for improvement for scientific tasks involving small molecules, proteins, DNA, RNA, or materials. In particular, foundation models struggle with precise quantitative predictions (e.g., ligand-protein binding affinity, protein-protein interactions, DNA properties) ai4science2023impactlargelanguagemodels and rational design and optimization of compounds, proteins, and materials. Moreover, ensuring the scientific accuracy of outputs from these models remains as a grand challenge.
@@
    Scientific data enhanced large language models (LLMs). This approach BioGPT2022Luo ; liu2023molxpt ; galactica2022 trains LLMs from scratch mainly with text data and a small portion of scientific data.
@@
The existence of these limitations emphasizes the necessity for a science foundation model, to fulfill the sophisticated demands of scientific research. A model of this kind must not only be highly proficient in producing precise scientific predictions but also adept at designing and optimizing scientific entities conditioned on context information. A perfect science foundation model ought to have the capacity to handle a diverse range of inputs. These inputs can span from literature text, to scientific sequence data such as protein or DNA sequences, and further to structural data like 3D protein/DNA structures and their dynamic behaviors. In the present study, our focus is on sequence-based data for representing biological, chemical, material systems, and natural language:
@@
We introduce Nature Language Model (briefly, NatureLM), a sequence-based science foundation model tailored for scientific tasks. NatureLM is designed to handle the complexity of small molecules, proteins, DNA, RNA, materials, and their associated textual information. NatureLM follows the Transformer decoder architecture and is trained on a corpus of 143 billion tokens collected from various scientific domains. Our experiments demonstrate that NatureLM significantly outperforms general-purpose foundation models for scientific tasks. Specifically, NatureLM excels in tasks such as:
@@
    Following textual instructions to generate/design and optimize scientific molecular entities.
@@
    Achieving state-of-the-art performance in retrosynthesis and SMILES-to-IUPAC translation.
@@
Figure 1: The overall ranking of models of varying sizes is depicted in the chart. The blue, orange, and green bars correspond to the results for the small (1B), medium (8B), and large (8x7B) models, respectively. A better rank is indicated by the “outsider” bar. The 8x7B model achieves top performance in 19 tasks, while the 8B model excels in 3 tasks. 18 categories exhibited performance improvements with increasing model size (i.e., 8x7B demonstrated the best performance, followed by 8B, and then 1B), highlighting the potential of large foundation models for scientific applications.
In summary, our development of the NatureLM represents a significant step towards building a generalist model across multiple scientific domains. By harnessing the capabilities of text-based instructions, NatureLM serves as a powerful tool for scientific discovery, enabling cross-domain generation and optimization in areas such as drug discovery, materials science, and the development of therapeutic proteins and nucleotides. Ideally, a foundation model should support a broad range of tasks while demonstrating strong zero-shot and few-shot capabilities. NatureLM shows great promise, but its language capabilities and few-shot learning skills still lag behind leading large language models. We will address these limitations in future iterations, positioning NatureLM as a key player in the continued evolution of scientific foundation models.
@@ 3.3 Translation between SMILES and IUPAC
We evaluate NatureLM on the translation between SMILES and IUPAC on NC-I2S and NC-S2I yu2024llasmol , the bidirectional IUPAC-SMILES translation dataset comprising 2993 pairs of SMILES and their corresponding IUPAC names (Table 4). We ensure that there is no test set leakage in this setting. On both text-to-SMILES and SMILES-to-text translation tasks, NatureLM (8x7B) outperforms all competing language models in terms of accuracy, demonstrating our model’s strong learning capability for text-molecule correspondence. NatureLM significantly outperforms GPT-4 and Claude 3 Opus claude3 , strong generalist large language models (LLMs), highlighting the necessity of training on scientific data. Compared with another LLM trained on text and SMILES corpus LlaSMolMistral yu2024llasmol , NatureLM also obtains significantly better performance. Moreover, NatureLM (8x7B) performs comparably with STOUT rajan2024stout , the widely-used model trained specially for IUPAC-SMILES translation task, demonstrating NatureLM’s potential as a scientific generalist in specific domains. The performance increases from NatureLM (1B) to NatureLM (8x7B), exhibiting the scaling benefits of larger models.
@@ 3.4 Target-aware hit generation and optimization
Vina QED SAS Diversity LogP$\in[0,5]$ Ro5
@@ 5.1 Unconditional material generation
We evaluated the SMACT validity of the generated materials. Furthermore, we used the dedicated fine-tuned NatureLM to autoregressively predict the crystal structures of a randomly chosen subset of valid compositions (see Section 5.4). The energy above hull (abbreviated as ehull) of the predicted structures was then evaluated using MatterSim yang2024mattersim . The distribution of ehull is shown in Fig. LABEL:fig:mat_uncon. We also assessed the ratio of stable materials, defining a generated material as stable if its ehull$<0.1$eV/atom. The results are presented in Table 12. It is evident that as the model size increases, the SMACT validity and stability of the generated materials improve.
@@ 5.3 Bulk modulus to material generation
Further, to demonstrate how NatureLM follows the instruction to generate materials with expected bulk modulus, we depict the distribution of the bulk modulus of generated materials under the instructions in Figure 18 where the $x$-axis denotes the bulk modulus in the instruction prompt and the $y$-axis denotes the bulk modulus of the generated materials under the corresponding instruction. We can see that, as the model scales, the distribution aligns more closely with the ideal linear diagonal.
@@
Figure 18: Distributions of bulk modulus of generated materials.
5.4 Crystal structure prediction for materials
@@
The fine-tuned NatureLM (1B) achieves performance that is comparable to or surpasses other state-of-the-art methods. The high match rates and low RMSE values demonstrate that our model effectively captures the complex spatial arrangements of atoms in crystal structures. Moreover we can see that NatureLM performs better than other methods as the number of atoms increases, demonstrating the advantage of autoregressive sequence model. As a next step, we plan to further improve the structure prediction quality by incorporating 3D autoregressive data into the pre-training phase of the next version of NatureLM.
@@ 9.1 Case study
Here we demonstrate our model’s language capabilities in scientific discovery research. Specifically, we assessed the model’s understanding of CRISPR gene editing, cancer immunotherapy, and drug discovery. A few selected cases are presented in this section for brevity. Additional cases can be found in Fig. S14, S15 and S16.
@@ 11.3 Cross-domain applications
Jianwei Zhu, Kehan Wu, Wei Zhang, Kaiyuan Gao, Qizhi Pei,
@@ Acknowledgements
We extend our gratitude to Dr. Fan Yang and Dr. Jilong Xue for their support with large-scale model training. Our thanks also go to Dr. Claudio Zeni, Dr. Robert Pinsler, Dr. Daniel Zügner, Dr. Andrew Fowler, Dr. Matthew Horton, and Dr. Ryota Tomioka for their assistance with material tasks. We appreciate the constructive feedback from Dr. Bichlien Nguyen, Dr. Jake Smith, and Dr. Frank Noé. We thank Dr. Han Guo for visualizing the molecules in our paper. We acknowledge Jingyun Bai for improving the quality of the figures. We thank Dr. Christopher Bishop for his invaluable guidance and sponsorship of this work.

### v2

# Nature Language Model: Deciphering the Language of Nature for Scientific Discovery

@@ Abstract
Foundation models have revolutionized natural language processing and artificial intelligence, significantly enhancing how machines comprehend and generate human languages. Inspired by the success of these foundation models, researchers have developed foundation models for individual scientific domains, including small molecules, materials, proteins, DNA, RNA and even cells. However, these models are typically trained in isolation, lacking the ability to integrate across different scientific domains. Recognizing that entities within these domains can all be represented as sequences, which together form the “language of nature”, we introduce Nature Language Model (NatureLM), a sequence-based science foundation model designed for scientific discovery. Pre-trained with data from multiple scientific domains, NatureLM offers a unified, versatile model that enables various applications including: (i) generating and optimizing small molecules, proteins, RNA, and materials using text instructions; (ii) cross-domain generation/design, such as protein-to-molecule and protein-to-RNA generation; and (iii) top performance across different domains, matching or surpassing state-of-the-art specialist models.
@@ 1 Introduction
Despite their success in general-purpose tasks, early investigations ai4science2023impactlargelanguagemodels highlight significant room for improvement in scientific tasks involving small molecules, proteins, DNA, RNA, or materials. In particular, foundation models struggle with precise quantitative predictions (e.g., ligand-protein binding affinity, protein-protein interactions, DNA properties) ai4science2023impactlargelanguagemodels , as well as the rational design of small molecule compounds, proteins, or materials. Ensuring the scientific accuracy of outputs from these models remains a grand challenge.
@@
    Scientific data enhanced large language models (LLMs). This approach, exemplified by works such as BioGPT2022Luo ; liu2023molxpt ; galactica2022 , trains LLMs from scratch mainly with text data and a small portion of scientific data.
@@
These limitations emphasize the necessity for a science foundation model, to fulfill the sophisticated demands of scientific research. A model of this kind must not only be highly proficient in producing precise scientific predictions, but also adept at designing and optimizing scientific entities conditioned on context information. A good science foundation model ought to have the capacity to handle a diverse range of inputs. These inputs can span from literature text, to scientific sequence data such as protein or DNA sequences, and further to structural data like 3D protein/DNA structures and their dynamic behaviors. In the present study, our focus is on sequence-based data for representing biological, chemical, material systems, and textual human language (e.g., English):
@@
Figure 1: NatureLM is a GPT-style generative model trained on a diverse range of data, including small molecule compounds, proteins, DNA, RNA, materials, and both general and scientific texts, amounting to a total of 143 billion tokens. It is built on existing large language models by integrating new vocabularies for scientific entities and jointly pre-training all components. After the pre-training, the model undergoes additional instruction tuning using millions of curated instructions from scientific fields. Options for reinforcement learning and dedicated fine-tuning are also available to boost performance on specific tasks. Users can engage with NatureLM through natural language inputs. The model excels in various domains, achieving top results in tasks such as retrosynthesis (Section 3.7), SMILES-to-IUPAC translation (Section 3.3), protein generation (Section 4.1) and material property prediction (Section 8.3), often matching or exceeding the capabilities of state-of-the-art specialized models.
NatureLM is designed to handle the complexity of small molecules, proteins, DNA, RNA, materials, and their associated textual information. An overview of NatureLM is in Fig. 1. NatureLM follows the Transformer decoder architecture and is trained on a corpus of 143 billion tokens collected from various scientific domains (Fig. 4). Our experiments demonstrate that NatureLM significantly outperforms general-purpose foundation models for scientific tasks. Specifically, NatureLM excels in tasks such as:
@@
    Following textual instructions to generate and optimize scientific molecular entities.
@@
    Achieving top performance on generation and translation tasks, such retrosynthesis (Section 3.7), SMILES-to-IUPAC translation (Section 3.3), protein generation (Section 4.1), matching or surpassing state-of-the-art specialist models.
@@
Figure 2: The scaling effect in NatureLM is obvious. The chart depicts the overall ranking of models with varying sizes, where a better rank is represented by the “outsider” bar. The 8x7B model achieves top performance in 19 tasks, while the 8B model excels in 3 tasks. 18 categories exhibited performance improvements with increasing model size (i.e., 8x7B demonstrated the best performance, followed by 8B, and then 1B), highlighting the potential of large foundation models for scientific applications.
In summary, NatureLM represents a significant step towards building a generalist model across multiple scientific domains. By harnessing the capabilities of text-based instructions, NatureLM serves as a powerful tool for scientific discovery, enabling cross-domain generation and optimization in areas such as drug discovery, materials science, and the development of therapeutic proteins and nucleotides. Ideally, a foundation model should support a broad range of tasks while demonstrating strong zero-shot and few-shot capabilities. NatureLM shows great promise, but its language capabilities and few-shot learning skills still lag behind leading large language models. We will address these limitations in future iterations, positioning NatureLM as an essential component in the continued evolution of science foundation models.
@@ 3.3 Translation between SMILES and IUPAC
We evaluate NatureLM on the translation between SMILES and IUPAC on NC-I2S and NC-S2I yu2024llasmol , the bidirectional IUPAC-SMILES translation dataset comprising 2993 pairs of SMILES and their corresponding IUPAC names (Table 4). We ensure that there is no test set leakage in this setting. On both text-to-SMILES and SMILES-to-text translation tasks, NatureLM (8x7B) outperforms all competing language models in terms of accuracy, demonstrating our model’s strong learning capability for text-molecule correspondence. NatureLM significantly outperforms GPT-4 and Claude 3 Opus claude3 , strong generalist large language models (LLMs), highlighting the necessity of training on scientific data. Compared with another LLM trained on text and SMILES corpus LlaSMolMistral yu2024llasmol , NatureLM also obtains significantly better performance. Moreover, NatureLM (8x7B) performs comparably with STOUT rajan2024stout , the widely-used model trained specially for IUPAC-SMILES translation task, demonstrating NatureLM’s potential as a scientific generalist in specific domains. The performance increases from NatureLM (1B) to NatureLM (8x7B), exhibiting the scaling benefits of larger models. A case study is presented in Fig. S11, comparing NatureLM with general large language models and highlighting the advantages of NatureLM in scientific tasks.
@@ 3.4 Target-aware hit generation and optimization
Vina ($\downarrow$) QED SAS Diversity LogP$\in[0,5]$ Ro5
@@ 3.7 Retrosynthesis prediction
Figure 10: Case study on retrosynthesis prediction. We evaluated the performance of NatureLM, DeepSeek-R1, and o3-mini-high for retrosynthesis prediction using a reaction from U.S. Patent ID US12018024B2 (not included in our training set), granted to Novartis on June 25, 2024. NatureLM successfully proposed the ground-truth reactants from the patent, while the outputs from the other two methods required further refinement to achieve the same level of accuracy.
We selected a reaction from a U.S. patent with ID US12018024B2 (granted to Novartis on June 25, 2024) as the case study. As shown in Figure 10, the product is a brominated heterocyclic compound. To synthesize such molecules, organic chemists typically begin by constructing the ring system, then followed by a halogenation reaction on the ring. The halogenated site is subsequently used in further syntheses, often coupling with other ring systems. In this context, our NatureLM model accurately predicted one of the most common brominating agent NBS (N-Bromosuccinimide) in this step, aligning perfectly with experimental results. DFT optimization calculations to reactant CCOC(=O)c1cc2ccnn2cn1 show that 3-position is the most nucleophilic. So bromination prefers at 3-position. This demonstrates NatureLM’s capability to effectively predict useful reactants in chemical reactions.
In contrast, DeepSeek-R1 deepseekai2025r1 model selected incorrect reactants. DeepSeek-R1 focused on constructing a heterocyclic ring system in this step. However, it incorrectly positioned the nitrogen and carbon atoms in the substrate, leading to the wrong outcome. Meanwhile, o3-mini-high model selected another correct route but not so convenient for the whole synthetic process. o3-mini-high might attempt to construct the acetyl side chain by transition-metal-catalyzed CO insertion reaction. Although radical mechanism can achieve this since iodo atom is more reactive than bromo atom, introducing the side chain at this stage may still lead to a side reaction at the bromo atom. In the actual synthesis, this acetyl side chain is generally constructed in previous steps. These cases illustrate that, so far, general LLMs cannot fully grasp the rationale behind chemical synthesis. Although they can generate some reasons and predictions, these do not align with common chemist strategies.
Another example is shown in Fig. S12, where NatureLM accurately predicts the reactants, while the outputs from DeepSeek-R1 and o3-mini-high require additional refinement.
@@ 4.4 Protein description generation
4.5 Heme-binding protein design driven by text and SMILES
NatureLM is capable of designing proteins to meet various requirements. In this case study, we demonstrate how to use NatureLM to design a protein that binds to heme cofactor.
Heme is an essential molecule, playing critical roles in oxygen transport and various enzymatic processes. Designing a protein that can bind to heme can contribute to advancements in therapeutic development, biochemical research, and synthetic biology.
The protein design practice is outlined as the following 4 steps:
Step 1: protein generation. We prompt NatureLM to generate proteins using two approaches:
  1. 1.
    Text-based protein design, where we provide a text instruction (here in English);
  2. 2.
    Small-molecule-based protein design, where we use the SMILES representation of heme as part of the input instruction.
Examples of these prompts and their corresponding responses are shown below:
Example 4.2.
[…]
Instruction: Design a heme-binding protein sequence.
[…]
Instruction: Generate a protein that can bind to
[…]
Step 2: description generation. For each generated protein, we utilize the protein-to-text functionality in NatureLM (Section 4.4), to obtain a description of the properties and potential functions of the generated protein.
Step 3: screen proteins through keyword matching. In this step, we use GPT-4o to generate a keyword list, called HemeList, containing characteristics associated with heme-binding proteins. For every protein description generated in Step 2, we check whether it contains keywords from HemeList. If a description matches these criteria, the corresponding protein is added to a list called HemeProtein.
Step 4: structure generation and validation. For each protein in HemeProtein, we use Protenix Protenix2025 to predict the complex structure of the generated proteins bound to heme. The predicted structures are then inspected to ensure that the proteins can form the critical interaction with heme for stable binding.
Figure 17: Two examples of proteins with plausibility of binding to heme. The yellow models represent the generated protein structures, while the blue models correspond to the reference structures retrieved using the built-in “blast protein” function in ChimeraX chimerax2023 . In each model, the heme binding region is highlighted by showing the nearby residues in stick representations. We use the protein-to-text functionality of NatureLM to generate functional annotations for these proteins, and the original outputs are provided here: (left protein) “Heme-binding protein”; (right protein) “Transfers electrons from cytochrome c551 to cytochrome oxidase; C-type cytochrome; Part of the cbb3-type cytochrome c oxidase complex.”
In the text-based design case study (Fig.17 left), two histidine residues are positioned in close proximity to the iron located in the center of heme, enabling the formation of coordinated bonding interactions with the heme group. Similarly, in the SMILES-based design, the algorithm can output proteins with binding motifs similar to those generated in the text-based example (Fig. S13). However, as shown in Fig. 17 (right), we show a representative case where a methionine and histidine residue are observed to interact closely with the iron ion (see Fig. S14 for more discussion on this case). These residues effectively coordinate the metal ion through their respective side chains, demonstrating alternative structural strategies for heme binding. Furthermore, the designed protein sequences differ significantly from those present in the database, indicating that our approach can generate novel sequences with distinct structural properties. Collectively, these results demonstrate the effectiveness of NatureLM in designing functional heme-binding proteins with diverse and novel structural features. We also compare the apo and holo structures of the generated proteins in Fig. S15, which shows that the key residues involved in heme binding, such as histidine and methionine, occupy similar positions in both structures.
@@ 5 Material tasks
Crystal structures are fundamental to determining the physical, chemical, and mechanical properties of materials. Complementing the sequence outputs of NatureLM, we finetune NatureLM into a specialized model, NatureLM-Mat3D, which predicts crystal structures from the generated chemical formulas (see Section 5.4). After generating material formulas and space groups with NatureLM, we then utilize NatureLM-Mat3D to convert them into crystal structures for further evaluation and practical application.
@@ 5.1 Unconditional material generation
We evaluated the SMACT validity of the generated materials. Furthermore, we utilized NatureLM-Mat3D to predict the crystal structures of a randomly selected subset of valid compositions. The energy above hull (abbreviated as ehull) of the predicted structures was then evaluated using MatterSim yang2024mattersim . The distribution of ehull is shown in Fig. LABEL:fig:mat_uncon. We also assessed the ratio of stable materials, defining a generated material as stable if its ehull$<0.1$eV/atom. The results are presented in Table 12. It is evident that as the model size increases, the SMACT validity and stability of the generated materials improve.
@@ 5.3 Bulk modulus to material generation
Further, to demonstrate how NatureLM follows the instruction to generate materials with expected bulk modulus, we depict the distribution of the bulk modulus of generated materials under the instructions in Figure 21 where the $x$-axis denotes the bulk modulus in the instruction prompt and the $y$-axis denotes the predicted bulk modulus of the generated materials calculated by MatterSim. We can see that, as the model scales, the distribution aligns more closely with the ideal linear diagonal.
@@
Figure 21: Distribution of predicted bulk modulus values for generated materials. The $x$-axis represents the input bulk modulus values from the instructions, while the $y$-axis shows the predicted values for the generated molecules calculated by MatterSim.
Figure 22: Two cases with bulk modulus values near 400 GPa (evaluated via DFT) were identified. The chemical formulas, space groups, energy above the hull (e_hull), and bulk modulus values obtained from MatterSim and DFT calculations are provided.
Materials with an ultra-high bulk modulus are highly sought after due to their exceptional stiffness and incompressibility, making them indispensable for applications in extreme environments, such as aerospace, industrial tooling, and advanced engineering. To evaluate the potential of NatureLM in generating materials with high bulk modulus, we conducted a detailed analysis of the generated compositions targeted at a bulk modulus of 400 GPa. From the generated outputs, we manually identified cases where MatterSim yang2024mattersim predicted bulk modulus values within a 5 GPa range of the target. Two such cases were selected for further validation using density functional theory (DFT) calculations (see Fig. 22). The DFT results revealed bulk modulus values of 390 GPa and 394 GPa, which closely align with the target value of 400 GPa.
Beyond achieving the bulk modulus target, the two generated structures were confirmed to be novel compared to those available in the Materials Project database. This novelty underscores NatureLM’s potential for discovering new materials with exceptional mechanical properties, thereby broadening the scope of material design and innovation.
5.4 NatureLM-Mat3D: a crystal structure predictor for materials
@@
NatureLM-Mat3D (1B) achieves performance that is comparable to or surpasses other state-of-the-art methods. The high match rates and low RMSE values demonstrate that our model effectively captures the complex spatial arrangements of atoms in crystal structures. Moreover we can see that NatureLM-Mat3D performs better than other methods as the number of atoms increases, demonstrating the advantage of autoregressive sequence model. As a next step, we plan to further improve the structure prediction quality by incorporating 3D autoregressive data into the pre-training phase of the next version of NatureLM.
@@ 9.1 Case study
Here we demonstrate our model’s language capabilities in scientific discovery research. Specifically, we assessed the model’s understanding of CRISPR gene editing, cancer immunotherapy, and drug discovery. A few selected cases are presented in this section for brevity. Additional cases can be found in Fig. S19, S20 and S21.
@@ 11.3 Cross-domain applications
Jianwei Zhu, Ran Bi, Kehan Wu, Wei Zhang, Kaiyuan Gao, Qizhi Pei,
@@ Acknowledgements
We extend our gratitude to Dr. Fan Yang and Dr. Jilong Xue for their support with large-scale model training. We thank Likun Dong and Junren Li for conducting the case study on retrosynthesis and the SMILES-to-IUPAC translation. Our thanks also go to Dr. Claudio Zeni, Dr. Robert Pinsler, Dr. Daniel Zügner, Dr. Andrew Fowler, Dr. Matthew Horton, and Dr. Ryota Tomioka for their assistance with material tasks. We appreciate the constructive feedback from Dr. Bichlien Nguyen, Dr. Jake Smith, and Dr. Frank Noé. We thank Dr. Han Guo for visualizing the molecules in our paper. We acknowledge Jingyun Bai for improving the quality of the figures. We thank Dr. Christopher Bishop for his invaluable guidance and sponsorship of this work.
@@ Appendix A Supplementary figures
Figure S11: We selected SMILES strings from PubChem with IDs 172655007 and 172655008, which were available as of February 24, 2025, and were excluded from our training set. The performance of NatureLM, DeepSeek-R1 deepseekai2025r1 , GPT-4o, GPT-4.5-preview, and o3-mini was evaluated for SMILES-to-IUPAC translation. The generated IUPAC names are presented in the accompanying figure. These IUPAC names were subsequently converted back to SMILES for validation. The IUPAC name produced by o3-mini could not be processed due to the high structural complexity of the corresponding molecule. NatureLM successfully generated the correct result. It is important to emphasize that our objective is not to criticize the limitations of general language models but to better understand their current capabilities and explore how they can be complemented by NatureLM for enhanced performance. The molecular structures were visualized using the ChemDB Chemoinformatics Portal Chen2007-il https://cdb.ics.uci.edu/cgibin/Smi2DepictWeb.py.
Figure S12: Additional examples on retrosynthesis prediction. We evaluated the performance of NatureLM, DeepSeek-R1, and o3-mini-high using a reaction from U.S. Patent ID US11999726B2, granted to Eli Lilly on June 04, 2024. The product features two ring systems with a protecting functional group, suggesting that the previous synthesis step likely involved a reaction to connect these rings. Notably, an ether bond links the two rings, with a pyrazine ring on one side and a piperidine ring on the other. Substitution on the pyrazine ring is a common strategy due to its electrophilicity, which often leads to substitution reactions. In this case, NatureLM accurately predicted the cleavage site of the molecule, incorporated a common chlorine atom on the pyrazine ring, and preserved the molecule’s stereochemistry, providing a reasonable synthetic strategy. In contrast, both DeepSeek-R1 and o3-mini-high models correctly identified the reactive sites but failed to predict the correct reactants due to poor handling of SMILES representations. For instance, DeepSeek-R1 predicted the pyrazine as pyrimidine, altering the nitrogen atom’s position, while o3-mini-high converted the six-membered pyrazine directly into a five-membered imidazole. These errors indicate that these general-purpose language models do not fully understand the relationship between chemical structures and their SMILES representations, hindering their ability to perform accurate reaction predictions.
Figure S13: Additional examples of designing heme-binding proteins based on text or SMILES instructions are shown. The first two rows display results from the text-based design, while the second row corresponds to the SMILES-based design. The yellow models represent structures generated by NatureLM, whereas the blue models are the reference structures retrieved using the built-in Chimera function. The structures of the generated proteins were predicted using Protenix Protenix2025 .
Figure S14: Comparison of the complex structure of the generated protein with heme (yellow model) and heme C (pink model). The protein was obtained using the SMILES-to-protein approach described in Section 4.5. We observe that they share common structural features. The structures of the generated proteins were predicted using Protenix Protenix2025 .
For the retrieved PDB structure 3MK7 in Fig. 17, our generated protein aligns to the pocket region that binds to heme C. To further validate this, we used Protenix to predict the binding of our generated protein to both heme C (PubChem CID: 11987638) and heme. The results demonstrate that heme C fits properly into the designed pocket, supporting the structural compatibility of the generated protein with heme C.
This discrepancy arises from the high structural similarity between heme and heme C, as their SMILES representations are nearly identical. Despite this slight misalignment, the output remains biologically relevant because heme-binding proteins often interact with multiple heme derivatives. Furthermore, generating a protein that binds to heme C from the SMILES of heme highlights the algorithm’s ability to capture the inherent structural flexibility and functional overlap within the heme family. We will continue improving the algorithm to enhance ligand specificity in future iterations.
Figure S15: Comparison of the apo structure of the generated protein, the holo structure in complex with heme, and their aligned structures. Key residues, such as histidine and methionine, occupy similar positions in the pocket region in both the apo and holo structures. This observation suggests that the generated proteins are not only capable of binding heme but also exhibit a structurally pre-formed or conserved binding pocket even in the absence of the ligand. These findings validate the structural plausibility of the designed proteins and their suitability for heme binding.